Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

49. It is known that the sum of all the probabilities in a probability distribution is one.

(i) Sum of the probabilities = 0.4 + 0.4 + 0.2 = 1

Therefore, the given table is a probability distribution of random variables.

(ii) It can be seen that for X = 3, P (X) = −0.1

It is known that proba

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

48. If A⊂B, then A∩B=A

P(A∩B)=P(A) also P(A)<P(B)

Consider,

P(A/B)=P(A∩B)P(B)=P(A)P(B)≠P(B)P(A)

⇒P(A/B)=P(A∩B)P(B)=P(A)P(B) _____(1)

We know, P(B)≤1 ⇒1P(B)≥1

From eq. (1) , we have

P(A)P(B)≥P(A)

⇒P(A/B)≥P(A) _____(2)

⇒P(A/B) is not less than P(A)

Hence, from eq. (2) it can be concluded that the relation given in alternative (C) is correct.

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

47. Let, E1 and E2 be the event such that

E1: A speak truth

E2: A speak false

X: that head appears

P(E1)=45 and P(E2)=1−P(E1) = =1−45=15

If a coin tossed, then it may result either head (H) or tail (T) . The probability of getting a head is 12 whether A&nbs

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

46. Let, E1: event of choosing a diamond card

E2: event of choosing a card which is not diamond

A: denote the lost card, out of 52 cards

We know,

13 cards are diamond

39 cards are not diamond

P(E1)=1352=14

P(E2)=3952=34

When one diamond card is lost, there are 12 diamond cards out of 51 card

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

45. Let, E1: event of time consume by A

E2: event of time consume by B

E3: event of time consume by C

P(E1)=50%=50100=12

P(E2)=30%=30100=310

P(E3)=20%=20100=15

Let, A: event of producing the defective item. Therefore,

P(A/E1)=1%=1100

P(A/E2)=5%=5100

P(A/E3)=7%=7100

Therefore, by Baye’s theorem,

P(E1/A)= probability that the defective item was produced by A ,

P(E1/A)=P(E1).P(A/E1)P(E1).P(A/E1)+P(E2).P(A/E2)+P(E3).P(A/E3)

=12×110012×1100+310×5100+15×7100

=12×11001100(12+1510+75)=12×11001100(5+15+1410)=123410

=12×1034=534

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

44. Let, E1: event that outcome on the die is 5 or 6

E2: event that outcome on the die is 1,2,3 or 4

A: event of getting exactly one head

P(E1)=26=13

P(E2)=46=23

P(A/E1)= probability of getting exactly one head by tossing the coin three times if she gets 5 or 6 =38

P(A/E2)= probabilit

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

43. Let, E1: event that first group will win

E2: event that second group will win

A: event that new product will produce

P(E1)=0.6=610=35

P(E2)=0.4=410=25

P(A/E1)=P (introducing new product by group A ) =0.7=710

P(A/E2)=P ( introducing new product by group B ) =0.3=310

Therefore, by Baye’s theorem,

P(E2/A)= probability that n

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

42. Let, E1: event items produced by A

E2: event items produced by B

X: event that produced item was found to be defective

P(E1)=60%=35

P(E2)=40%=25

P(X/E1)=P (items produced by machine A which is defective) =2%=2100

P(X/E2)=P ( items produced by machine B which is defective) =1%=1100

Therefore, by Baye’s

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

41. Let, E1: event that the driver is scooter driver

E2: event that the driver is car driver

E3: event that the driver is truck driver

A: event the person meets with accident

Total number of drivers =2000+4000+6000=12000

P(E1)=P (driver is a scooter driver) =200012000=16

P(E2)=P (driver is a car driver) =400012000=13

P(E3)=P (dri

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

40. Let, E1: event of choosing Q headed coin

E2: event of choosing biased coin

E3: event of choosing unbiased coin

P(E1)=P(E2)=P(E3)=13

Let A be the event that shows head,

P(A/E1)=P (coin shows heads, given that it is a headed coin) =1

P(A/E2)=P (coin showing up head, given that it is biased coin

...Read more

Thumbs Up IconUpvote Thumbs Down Icon